HETAV PATEL

Grade 11,

prove that cos(π+teta)cos(-teta) / sin(π- teta)cos(π/2+ teta) = cot²q

prove that cos(π+teta)cos(-teta) / sin(π- teta)cos(π/2+ teta) = cot²q

Grade:12th pass

1 Answers

Arun
6 years ago
we can write
cos(\pi\pie + theta) =  –cos (theta)
cos( – theta) = cos (theta)
sin(\pi- theta) = sin (theta)
cos(\pi/2 + theta) =  –sin(theta)
now put these value in L.H.S.
cos2(theta) / sin2(theta) = cot2(theta)
hope it clears